Jim aged 50 years, has a daughter Diana of age 20 years. How many years ago was the age of Jim three times that of his daughter?
Answer by Maths68(1474)

Present age of Jim = 50
Present age of Diana = 20
Let
x years ago Jim was three times that of his daughter


x years ago Jim's age = 50-x
x years ago Diana's age = 20-x
x years ago the age of Jim was three times that of his daughter
50-x=3(20-x)
50-x=60-3x
3x-x=60-50
2x=10
x=10/2
x=5
Fiver years ago Jim's age was three times that of Diana's
Diana was 15 and Jim was 45 (which is 3 times of 15)